# Relative pose of three calibrated and partially calibrated cameras from four points using virtual correspondences

### v2

@@ 4 Experiments
We evaluated the proposed solvers on synthetic and real data to test their robustness to noise and outliers, and assess their performance inside a state-of-the-art RANSAC-framework [3]. We compare our novel solvers with the following state-of-the-art algorithms:
  • [leftmargin=0.3cm]
  • consecutive 5pt/6pt: RANSAC-based robust relative pose estimation is applied independently for two pairs of views (1-2 and 1-3), without considering triplets.
  • joint 5pt/6pt: inside a single RANSAC, from a sample of 5/6 points in 3 views, essential matrices (and the focal length) are estimated independently for two pairs of views (1-2 and 1-3) from the view triplet. The two matrices are jointly validated inside RANSAC.
  • 5pt+P3P/6pt+P3P: The baseline minimal solvers for the $(5_{3},5_{3},3_{3})$ and the $(6_{3},6_{3},3_{3})$ configurations for calibrated/cameras with unknown focal length.
  • 4p3v(HC): The homotopy continuation solver [21].
[…]
For the synthetic experiment, we generated 10k scenes with known ground truth parameters. In each scene, the three 3D points were randomly distributed within a cube of size $10\times 10\times 10$. Each 3D point was projected into two cameras. The orientations and positions of the cameras were selected at random such that they looked towards the origin from a random distance, varying from $20$ to $50$, from the scene. In the first camera, we generated a new point as a mean point of the three projected points. In the second camera, we were sampling points inside the triangle defined by the three 2D points. To make the sampling invariant of actual coordinates of 2D points, we represented sampling in barycentric coordinates. We uniformly sampled $19\times 19$ barycentric coordinates $(a,b)\in[0,1]^{2}$, such that $a+b\leq 1$. The third barycentric coordinate was computed as $c=1-a-b$. We use these barycentric coordinates to compute the coordinates of the corresponding point in the second image. We then evaluated the symmetric epipolar error of this point and the mean point in the first image. Fig. 2 (left) shows the average symmetric epipolar error as a function of the barycentric coordinates of the point in the second image. The 2D Gaussian distribution fitted to these results has mean $\mu=(0.333,0.332)$. In this point the minimum was reached. The point with barycentric coordinates $(0.\bar{3},0.\bar{3})$ corresponds to the mean point of the triangle.
Figure 2: Distribution of the average symmetric epipolar error as a function of barycentric coordinates of the point in the second image w.r.t. the mean point of three points in the first image, for synthetic data (left) and real data (right).
(a) 0.1 Inlier Ratio
(b) 0.2 Inlier Ratio
(c) 0.4 Inlier Ratio
Figure 3: Outlier Experiments: We tested the performance, in terms of average pose error (maximum of rotation and translation errors) of the solvers for different inlier ratios and for different numbers of RANSAC iterations. The result is for the Sacre Coeur scene from Phototourism dataset [54]. We plot the results for (a) 0.1 inlier ratio, (b) 0.2 inlier ration and (c) 0.4 inlier ratio.
We performed the same experiment on real data. In this experiment we used the SfM model of the Shop Facade scene from the Cambridge Landmark dataset [23] to obtain image pairs. We extracted 4,525 pairs and used exact projections of 3D points into these pairs to simulate noise-less data. In each pair, we sampled 100 different triplets of points and for each, we performed the same experiment as described above. Altogether, we evaluated 452,500 triplets (corresponding triangles). Fig. 2 (right) again shows the average symmetric epipolar error as a function of barycentric coordinates of the point in the second image. The distribution of errors was very similar to the synthetic data with the minimum at $(0.333,0.333)$, i.e., the mean point.
Noise experiments.  We tested the performance of our solvers, and the state-of-the-art algorithms w.r.t. increasing image noise. We use the SfM model of the botanical garden scene from the ETH3D dataset [53] to obtain instances of points in three views. Perfect correspondences with no noise are generated by projecting the 3D points into the images. We then add increasing amounts of normally distributed noise to these correspondences. Due to the space restrictions, the results of this experiment are presented in the supp. material.
Here we just summarize the main observations from this experiment: Due to the approximate nature of virtual correspondences, our novel solvers return non-zero errors for zero noise. However, at noise levels $\geq 2px$, these solvers return comparable or even better results than the 5pt+P3P/6pt+P3P solvers. This shows that our predicted virtual correspondences are good approximations of real correspondences. The recent state-of-the-art solver [21] is failing in about 50% of the instances for noiseless data, even though the solver was trained on the ETH3D dataset. Thus, the median errors are significantly larger than the median errors of the remaining solvers.
Oracle 4v3p solvers.  To obtain upper bounds for the precision that can be achieved by our proposed solvers, as well as the state-of-the art 4p3v(N) solver [39], we consider the following “oracle” solvers for real experiments: 1) The 4p3v(O)/4p3vf(O) solvers use a correct correspondence(s), i.e., correspondences that satisfies the epipolar constraint for the ground truth relative pose of the first two cameras, as the $5^{\text{th}}$/$6^{\text{th}}$ virtual correspondence between these cameras. Then the 5pt+P3P/6pt+P3P solver is applied to estimate the relative pose of three cameras. The 4p3v(O)/4p3vf(O) shows the maximum precision that our solvers can reach, if they would have been able to predict or infer a precise $5^{\text{th}}$/$6^{\text{th}}$ correspondence from the coordinates of four input correspondences. 2) We test an oracle version of the 4p3v(N) solver. In this 4p3v(NO) solver, instead of performing a one-dimensional search over the $10^{\text{th}}$ degree curve of possible epipoles, we provide the solver with the ground truth epipole. More details on this solver are in the supp. material.
Sacre Coeur (5000 image triplets) St. Peters’ Square (5000 image triplets)
Estimator AVG (${}^{\circ}$) $\downarrow$ MED (${}^{\circ}$) $\downarrow$ ${}^{\circ}$uparrow ${}^{\circ}$${}^{\circ}$ Time (s) AVG (${}^{\circ}$) $\downarrow$ MED (${}^{\circ}$) $\downarrow$ ${}^{\circ}$uparrow ${}^{\circ}$${}^{\circ}$ Time (s) Time ($\mu$s)
cons. 5pt 14.24 15.06 35.20 45.82 56.76 10.15 16.46 19.12 16.49 30.27 46.39 10.04 1113.38
joint 5pt 24.88 14.33 38.45 47.32 55.25 10.29 26.36 10.47 15.31 28.15 43.38 10.36 1113.38
5pt+P3P 23.80 11.86 47.69 55.13 61.25 13.32 26.02 18.67 20.85 33.69 47.17 13.13 1177.90
4p3v(HC) [21] 24.72 13.04 42.81 50.61 57.27 12.60 27.06 19.73 19.78 32.03 45.10 12.22 1*66.06
4p3v(M) 24.77 13.08 42.62 50.58 57.52 13.12 27.44 19.88 18.59 30.99 44.61 12.74 1183.92
4p3v(M $\pm$ 0.15) 24.49 12.42 44.51 52.27 58.91 13.10 25.96 18.92 19.62 32.53 46.24 12.76 1218.71
4p3v(L) 24.92 12.97 42.84 50.71 57.51 13.82 26.78 19.49 18.73 31.51 45.14 12.93 1450.26
4p3v(L $\pm$ 0.1) 24.36 12.35 44.89 52.48 59.00 13.76 26.12 19.06 19.77 32.52 46.06 13.39 1511.28
4p3v(L $\pm$ 0.1 init) 23.95 12.46 44.83 52.78 59.29 15.10 26.09 18.83 19.67 32.68 46.48 14.67 1130.31
4p3v(O) 21.91 11.54 50.18 57.68 63.62 12.85 24.91 16.45 25.67 39.04 51.80 13.11 1180.47
4p3v(NO) 23.34 12.13 46.15 54.54 61.08 10.22 25.89 17.90 21.23 34.79 48.34 10.20 1*73.65
Table 1: The average and median pose errors in degrees, and Area Under the Recall curve (AUC) thresholded at ${}^{\circ}$, ${}^{\circ}$ and ${}^{\circ}$ and the average run-time of RANSAC (in seconds) on a total of 10k image triplets from two scenes of the PhotoTourism dataset [22]. The best results among the solvers for the 4p3v problem (excluding oracle solvers) are marked in bold, the second best are underlined. The last column contains the average running times of particular solvers in $\mu$s over 1,000 trials. $*{}^{*}$ The time for 4p3v(HC) does not account for its high failure rate, the effective time is usually more than 4×4$\times$4 × higher [21]. Our time for the 4p3v(HC) is higher than the average time reported in [21]. This can be due to different instances and hardware, and that time in [21] did not involve, e.g., pose extraction from estimated depths. The time for 4p3v(NO) is the time to evaluate one search point. Usually 40 – 1,000 search points have to be evaluated.
Timings.  We computed the average processing times of the solvers over 1,000 random problem instances. The solvers were implemented in C++ within the GC-RANSAC [4] framework using the Eigen library for performing matrix operations. We used the 5pt [38], 6pt [19] and P3P [43] methods implemented in GC-RANSAC for composing the discussed solvers. For 4p3v(HC) we used the implementation from the authors [^5] https://github.com/petrhruby97/learning_minimal with default settings [21]. The experiments were performed on an Intel(R) Core(TM) i9-10900X CPU @ 3.70GHz. The results for calibrated solvers are reported in the last column of Tab. 1. The discussion of these results is included in Tab. 1 and in Sec. 5. Timings for solvers for partially calibrated cameras are in the supp. mat.
Outlier experiments.  The advantage of 4p3v/4p3vf solvers for the $(4_{4},4_{4},4_{4})$ configuration over the 5pt+P3P and 6pt+P3P solvers for the $(5_{3},5_{3},3_{3})$/$(6_{3},6_{3},3_{3})$ configuration should be visible, especially in the presence of higher outlier contamination and a limited running time for RANSAC. To study this scenario, we evaluate the performance of all proposed solvers, the state-of-the-art 4p3v(HC) solver [21], and the 5pt+P3P/6pt+P3P solvers w.r.t. decreasing inlier ratio, while reporting errors and metrics on different number of RANSAC iterations. For calibrated case we use the Sacre Coeur and St. Peters’ Square scenes from PhotoTourism dataset [54]. We run the solvers inside RANSAC [4] and measure their performance on a series of iterations, only on triplets of images with a given inlier ratio. We used the uniform sampler inside RANSAC. To obtain triplets of images with the given inlier ratio, for each triplet, we keep a certain amount of inliers depending on the total number of correspondences, and replace the rest by uniformly random outlier w.r.t. ground truth poses, correspondences. We filter out image triplets that do not contain enough inliers. The results of this experiment for calibrated cameras and inlier ratios $\{0.1,0.2,0.4\}$ are presented in Figure 3. The results indicate that the new proposed 4p3v solvers reach lower errors faster than the 5pt+P3P solver that samples one more point in the first two cameras. The HC [21] solver, due to its high failure rate, converges even slower than the 5pt+P3P solver.
Real data experiments.  We tested the solvers on two scenes from the validation set of the IMC2020 benchmark [22]. The scenes, Sacre Coeur and St. Peter’s Square, come from the PhotoTourism dataset [54] and provide ground truth pose via a COLMAP [52] reconstruction. To form image triplets, we iterated through the point correspondences provided in [22] and selected pairs of image pairs that share one view. Since the provided correspondences do not contain descriptors, we detected RootSIFT [35, 2] features in all three images of each triplet. We then matched features between the view pairs 1-2 and 1-3 by standard nearest neighbors matching. To enforce cycle consistency, we kept only those correspondences that were matched to the same features in all view pair combinations. In total, we collected 5,000 image triplets from each scene.
To estimate the relative poses between the images of each triplet, we apply the Graph-Cut RANSAC [4] robust estimator. The point-to-model residual function is chosen to be the Sampson error of a triplet correspondence average over image pairs 1-2 and 1-3. We calculate the error of the triplet pose by, first, taking the average of the relative rotation and translation errors coming from essential matrices $\textbf{E}_{12}$ and $\textbf{E}_{13}$ given the ground truth relative pose. Finally, we take the maximum of the averaged rotation and translation errors (both measured in degrees) as the pose error [22].
We test the focal length solvers on the KITTI dataset [17]. We form triplets by connecting the stereo pair from the $i$-th frame to the left image in the $i+1$-th one. Similarly as before, we extract and match RootSIFT features.
Tab. 1 and 2 report the average and median pose error in degrees, the Area Under the recall Curve (AUC) thresholded at $5^{\circ}$, $10^{\circ}$, and $20^{\circ}$, and the average processing time in seconds. The results are discussed in Sec. 5. More results on real data are in the supp. material.
5 Discussion and limitations
In this paper, we present solvers to the challenging problems of estimating the relative pose of three calibrated and partially calibrated cameras from four correspondences. We tested our solvers on different real-world datasets (Tabs. 1 and 2) from which several conclusions can be drawn:
  • The very simple 4p3v(M±$\pm\delta$)/4p3vf(M±$\pm\delta$) solvers, which use only mean points as virtual “approximate” correspondences and two more points in their vicinity, achieve state-of-the-art results on the challenging 4p3v/4p3vf problems. Compared to state-of-the-art solvers for the 4p3v problem, our new 4p3v(M±$\pm\delta$)/4p3vf(M±$\pm\delta$) solvers are extremely simple to implement since they just run existing well-known 5pt/6pt and P3P solvers that are available in many libraries. They do not require any network training, parameter tuning, optimization, or implementation of complex solution techniques as required by the 4p3v(HC) and 4p3v(N) solvers. As such, our novel M-solvers should be considered as new baselines for the 4p3v/4p3vf problems.
  • The 4p3v(M±$\pm\delta$)/4p3vf(M±$\pm\delta$) solvers call the 5pt/6pt solvers three times and are thus slower than the 4p3v(M)/4p3vf(M) solvers. However, since $\delta$-solvers provide better estimates, the total running times of RANSAC with the 4p3v(M±$\pm\delta$)/4p3vf(M±$\pm\delta$) solvers are almost identical to the running times of RANSAC with the 4p3v(M)/4p3vf(M) solvers, and comparable to the 4p3v(HC) [21], while achieving better precision.
  • Our novel L-based solvers for the 4p3v problem show a slight improvement over the M-based solvers (cf. Tab 1). Although the architecture of our proposed network is very simple and can be considered more as a prototype, it shows the potential to learn virtual correspondences and achieve results closer to the oracle 4p3v(O) solver.
  • For the 4p3vf problem with unknown focal length, the proposed 4p3vf(M±$\pm\delta$) solver achieves results very close to the oracle 4p3vf(O) solver (cf. Tab. 2). For this problem, the 4p3vf(L) solver achieves smaller average but larger median errors than the 4p3vf(M) solver. We attribute this to the fact that we use a network trained for the 5pt solver, while the 6pt solver exhibits more degenerate configurations, and the fact that for this dataset, mean point correspondences are most likely very good approximations of correct correspondences.
  • The state-of-the-art results of our solvers show the potential of the proposed method based on virtual correspondences to be applied to other camera geometry problems. A discussion on this can be found in the supp. material.
  • We are the first to properly compare solvers for two different minimal configurations, i.e., the $(5_{3},5_{3},5_{3})$ and $(4_{4},4_{4},4_{4})$ for three calibrated cameras on real-world datasets within a state-of-the-art RANSAC framework [4]. Although our novel solvers achieve state-of-the-art results for the $(4_{4},4_{4},4_{4})$ configuration, our results show that in scenarios where we have enough computational time and a small outlier contamination, it is preferable to use the solver 5pt+P3P and sample one more point in two views. However, the results of the oracle 4p3v(O) solver suggest that there is still room for improving 4p3v solvers. Moreover, our outlier experiments (cf. Fig 3) show that in the presence of larger outlier ratios and limited computational time, the proposed 4p3v solvers are preferable.
Estimator AVG (${}^{\circ}$) $\downarrow$ MED (${}^{\circ}$) $\downarrow$ ${}^{\circ}$uparrow ${}^{\circ}$${}^{\circ}$${}^{\circ}$ Time (s)
cons. 6pt 11.69 11.14 11.77 47.34 70.30 84.57 10.05
joint 6pt 11.59 11.09 12.85 48.77 71.05 85.00 10.05
6pt+p3p 11.83 11.00 15.79 51.14 71.73 84.89 10.11
4p3vf(M) 11.89 11.05 14.47 49.64 70.81 84.44 10.07
4p3vf(M $\pm$ 0.15) 11.66 10.99 15.74 51.36 72.05 85.27 10.07
4p3vf(L) 11.58 11.10 12.88 48.79 71.09 85.06 10.07
4p3vf(L $\pm$ 0.1) 11.69 11.01 15.44 50.68 71.55 85.07 10.10
4p3vf(L $\pm$ 0.1 init) 11.74 11.01 15.30 50.30 71.34 84.84 10.18
4p3vf(O) 11.58 10.98 16.05 51.59 72.25 85.51 10.09
Table 2: Results on KITTI dataset [17] for partially calibrated cameras. The reported statistics are the same as in Tab. 1.
6 Conclusion
In this paper, we consider the highly challenging problem of relative pose estimation of three calibrated and partially calibrated cameras from four correspondences. We propose a novel framework that solves these problems using existing efficient solvers by simply predicting a $5^{\text{th}}$/$6^{\text{th}}$ point correspondence(s). We propose several solvers based on this framework, one simply using mean coordinates of input points (M-based solvers) and one using a trained predictor (L-based solvers). Extensive experiments show that our solvers achieve state-of-the-art performance on real data for the configuration of four points in three views. At the same time, our M-based solvers are trivial to implement, especially compared to the current state-of-the-art [21].
Acknowledgements.  Charalambos Tzamos was supported by the Czech Science Foundation (GAČR) JUNIOR STAR Grant (No. 22-23183M) and by SGS project (No. SGS23/173/OHK3/3T/13). Daniel Barath was supported by the ETH postdoc fellowship. Torsten Sattler was supported by the EU Horizon 2020 project RICAIP (grant agreement No. 857306) and the European Regional Development Fund under project IMPACT (No. CZ.02.1.01/0.0/0.0/15_003/0000468). Zuzana Kukelova was supported by the Czech Science Foundation (GAČR) JUNIOR STAR Grant (No. 22-23183M).
Supplementary Material
This supplementary material provides more details on the architecture, training data, and training of the network used in the 4p3v(L)/4p3vf(L) solvers presented in Sec. 3.1 and 3.2 of the main paper (Sec. 7), a discussion of the proposed method for solving camera geometry problems using virtual correspondences (Sec. 8), more details on the experiments presented in Sec. 4 of the main paper (Sec. 9), and potential directions for future work (Sec. 10).
7 Details on the network
As described in the main paper, we use a simple MLP based backbone, similar to [10]. More precisely, our input consists of four 6-vectors, of the $x,$ and $y$ coordinates of four point correspondences in three views. The first part of the model is a shared MLP 3-block of dimensions 6, 32, and 32, exporting a 32-dimensional feature for every correspondence. Then we apply a channel-wise max pooling aggregation, which is the concatenated at the end of each 32-dimensional feature. This results in having an 64-vector for each correspondence, which are passed into another shared MLP 3-block of 64, 64, and 64 dimensions. We aggregate those vector via a max pooling function to get a 64-vector encoding, which eventually passes through an MLP 3-block to reduce the dimension gradually from 64, to 32 and finally to 2, which will be the prediction of the $x$ and $y$ coordinates in the second camera. In all MLPs, in the first 2 layers of the blocks, we use a Leaky RELU activation function [62] with slope 0.01. As for the last layer of the MLPs, in the first two blocks, we use a RELU activation function, while in the last MLP we use a $\tanh$ activation, since we want the output to be in the range $(-1,1)$. For a visualization of the architecture, see Figure 5.
For both training and validation, we use synthetic data. Our synthetic dataset contains 1M input instances. 70% of the dataset is used for training while the rest 30% is used for validating the performance of the network. We generate 10K 3D points inside a 10$\times$10$\times$10 cube, and to generate each instance, we pick 4 random points and project them to 3 cameras with random rotations and translations that view the cube from a random distance between 20 to 50 units.
8 Solving camera geometry problems using virtual correspondences
The proposed methods generating virtual correspondences, i.e., the learning-based and the mean-point-based methods that were used in the novel solvers for the 4p3v and 4p3vf problems, can be, in general, applied to other camera geometry estimation problems.
M-based solvers  While the mean point correspondence used in the M-based solvers can provide a good approximation of a correct correspondence, such a correspondence can be noisy. Note that all state-of-the-art 4p3v solvers (including 4p3v(HC) and 4p3v(N)) rely on certain approximations without establishing theoretical proofs to quantify their accuracy. In contrast, the error of our virtual correspondence is bounded: As mentioned in the main paper, it can be proven that the error of the virtual correspondence $\mathbf{m}^{1}\leftrightarrow\mathbf{m}^{2}$ is bounded by the maximum distance of the mean point $\mathbf{m}^{2}$ from the vertices of the triangle $\left\{\mathbf{x}_{i}^{2},\mathbf{x}_{j}^{2},\mathbf{x}_{k}^{2}\right\}$. Here we provide a simple proof.
Figure 4: Illustration of the considered situation.
Lemma 1.
Let us assume two cameras with camera centers $\mathbf{C}^{1}$ and $\mathbf{C}^{2}$ that observe 3D points $X_{i},X_{j},$ and $X_{k}$. Let $\left\{\mathbf{x}_{i}^{1},\mathbf{x}_{j}^{1},\mathbf{x}_{k}^{1}\right\}$ and $\left\{\mathbf{x}_{i}^{2},\mathbf{x}_{j}^{2},\mathbf{x}_{k}^{2}\right\}$ be the projections of these 3D points in camera 1 and camera 2, respectively. Let $\mathbf{m}^{1}$ be the mean point of the points $\left\{\mathbf{x}_{i}^{1},\mathbf{x}_{j}^{1},\mathbf{x}_{k}^{1}\right\}$ and let $\mathbf{E}$ be the essential matrix between these two cameras, i.e., a matrix that satisfies $\mathbf{x}_{l}^{2^{\top}}\mathbf{E}\mathbf{x}_{l}^{1}=0,\;\forall l\in\left\{i,j,k\right\}$. Then the epipolar line $\mathbf{E}\mathbf{m}^{1}$ passes through the triangle $\left\{\mathbf{x}_{i}^{2},\mathbf{x}_{j}^{2},\mathbf{x}_{k}^{2}\right\}$.
Proof.
The camera center $\mathbf{C}^{1}$ and the 3D points $\mathbf{X}_{i},\mathbf{X}_{j},$ and $\mathbf{X}_{k}$ form a tetrahedron $T^{1}$ (see Figure 4.). The projections $\left\{\mathbf{x}_{i}^{1},\mathbf{x}_{j}^{1},\mathbf{x}_{k}^{1}\right\}$ in the first camera lie at the edges of this tetrahedron $T^{1}$. The ray from the camera center $\mathbf{C}^{1}$ through the mean point $\mathbf{m}^{1}$ thus lies inside the tetrahedron $T^{1}$ and intersects the plane defined by 3D points $\mathbf{X}_{i},\mathbf{X}_{j},$ and $\mathbf{X}_{k}$ in a point $\mathbf{M}$ that lies inside the triangle defined by $\left\{\mathbf{X}_{i},\mathbf{X}_{j},\mathbf{X}_{k}\right\}$.
Obviously, it follows from Lemma 1 that since the epipolar line $\mathbf{E}\mathbf{m}^{1}$ passes through the triangle $\left\{\mathbf{x}_{i}^{2},\mathbf{x}_{j}^{2},\mathbf{x}_{k}^{2}\right\}$, the maximum distance of the mean point $\mathbf{m}^{2}$ to the epipolar line $\mathbf{E}\mathbf{m}^{1}$ is equal to the maximum distance of $\mathbf{m}^{2}$ to the vertices of the triangle.
The final error of the estimated pose, however, depends on many aspects, e.g., the baseline and the view angles of the cameras w.r.t. the 3 points used, the depth of these points, and the size of the triangle. Although in theory, this error could be large, our experiments show that it does not lead to significantly larger errors than the 5pt+P3P method, which samples an additional fifth correspondence, rather than predicting or approximating one (e.g., for the ShopFacade from the Cambridge Landmarks dataset, avg/med epipolar error of the mean point prediction is 6.01px/2.94px).
A deeper analysis of the error propagation for virtual correspondences is out of the scope of this paper and better fits a follow-up publication dealing with the proposed method as a general method for solving camera geometry problems. In this paper, we decided to focus on two important problems of estimating the relative pose of three cameras from four point correspondences, which have been considered as not fully solved for decades.
L-based and $\delta$-based solvers  Our simple L-based and $\delta$-based solvers, that either predict “a shift” of the mean point in the second image using a simple network or generate two more potential correspondences in the $\delta$-vicinity of the mean-point correspondence, show the potential to deal with noise in the mean-point correspondences. Note that we tested only a simple network architecture and a simple strategy to generate point correspondences in the $\delta$-vicinity of the mean-point correspondence. More advanced network architectures (e.g., using a set transformer architecture [32]) or a more advanced method to generate additional $\delta$-point correspondences have the potential to improve the results even more and reduce the gap between our proposed 4p3v/4p3vf solvers and the oracle 4p3v(O)/4p3vf(O) solvers.
Use-cases for virtual correspondences  We believe that the proposed method based on virtual correspondences is general and, due to its simplicity, has the potential to generate efficient and easy-to-implement solvers for other problems. Generating one or potentially more virtual correspondences using only the locations of the input correspondences can allow to solve complex minimal problems using much simpler non-minimal solvers, while still sampling only a minimal number of points inside RANSAC.
In the first step, our M-based solvers solve 5pt/6pt relative pose problems of two calibrated/partially calibrated cameras using only a sub-minimal sample of four point correspondences. The good performance of these solvers, not only for the pose of the whole triplet of cameras but also for the relative pose of the first two cameras (cf. Sec. 9.2 and Figure 7), suggests that the method based on virtual correspondences can also be used to solve minimal problems using sub-minimal samples inside RANSAC in some scenarios.
To better illustrate this use-case, in Sec. 9.6 we show preliminary results for the well-known 5pt relative pose problem, where we sample only four point correspondences inside RANSAC and generate the $5^{\text{th}}$ correspondence as the mean point (or points in its $\delta$-vicinity) of three corresponding points in both images. Even though four points in two views define an infinite number of camera poses that can observe these points, we show that, in practice, such a mean-point correspondence is usually a quite good approximation of a correct correspondence. While this $5^{\text{th}}$ correspondence can be noisy, the resulting 4pt(M) and 4pt(M±$\pm\delta$) solvers show comparable performance to the 5pt solver inside RANSAC. This shows the potential for applying such a sub-minimal solver in scenarios with lots of outliers and noisy data, where sampling one less point may show some improvement.
While the method based on virtual correspondences can be theoretically applied to any camera geometry problem, we see larger promise in relative pose problems, where it is sufficient to find one 2D point that is sufficiently close to the epipolar line. For absolute pose solvers, a virtual correspondence will need to be close to a 2D point instead of an epipolar line.
9 Experiments
This section provides more details on the experiments presented in Sec. 4 of the main paper, as well as results showing the potential of the proposed method based on virtual correspondences to solve minimal problems using sub-minimal samples of points.
[…]

### v3

@@ 9 Experiments
This section provides more details on the experiments presented in Sec. 4 of the main paper. More precisely, Sec. 9.7 provides details and experiments on the oracle solvers discussed in the main paper. Sec. 9.1 provides ablation studies for the number of iterations of the refinement strategy (+R) and for the choice of $\delta$. Sec. 9.2 provides details on the noise experiments summarized in the main paper. Sec. 9.3 provides experiments with an evaluation measure taking the relative pose error between the 2nd and 3rd camera in a triplet into account. Sec. 9.4 presents results on the PhotoTourism dataset obtained with GC-RANSAC. Sec. 9.5 presents additional experiments with adding outliers to image triplets (similar to the experiments presented in Fig. 4 in the main paper). Sec. 9.6 presents the additional details on some experiments promised in Sec. 4 of the main paper. Finally, Sec. 9.8 measures the run-times of the different solvers considered in this work.
9.1 Ablation studies
Estimator $\delta$ AVG $(^{\circ})$downarrow MED $(^{\circ})$downarrow AUC@5 $\uparrow$ @10 $\uparrow$ @20 $\uparrow$
4p3v(M±$\pm\delta$) 0.2 6.33 3.89 36.88 57.46 74.74
0.1 6.28 3.82 37.50 57.88 74.95
0.09 6.29 3.77 37.98 58.27 75.18
0.08 6.31 3.72 38.32 58.53 75.36
0.07 6.21 3.64 39.29 59.16 75.69
0.06 6.18 3.68 39.16 59.18 75.76
0.05 6.07 3.65 39.47 59.37 75.93
0.04 5.99 3.62 39.57 59.52 76.01
0.03 6.04 3.64 39.44 59.36 75.92
0.02 6.18 3.68 38.74 58.73 75.42
0.01 6.30 3.81 37.89 57.94 74.90
0.005 6.39 3.88 36.62 56.94 74.36
0.001 6.65 4.03 35.57 55.83 73.50
4p3v(M±$\pm\delta$)+R 0.2 5.89 3.40 41.40 60.73 76.71
0.1 5.80 3.32 42.48 61.72 77.28
0.09 5.71 3.31 42.37 61.62 77.32
0.08 5.68 3.29 42.86 61.95 77.42
0.07 5.65 3.25 43.25 62.27 77.68
0.06 5.57 3.28 43.12 62.07 77.56
0.05 5.59 3.24 43.23 62.18 77.66
0.04 5.64 3.19 43.60 62.43 77.66
0.03 5.66 3.22 43.68 62.49 77.82
0.02 5.65 3.21 43.24 62.11 77.56
0.01 5.79 3.29 42.54 61.75 77.18
0.005 5.90 3.38 42.03 61.24 76.91
0.001 6.09 3.46 40.80 59.99 75.98
4p3v(M±$\pm\delta$)+R+F 0.2 6.02 3.60 39.44 59.19 75.79
0.1 6.00 3.50 40.69 60.27 76.36
0.09 5.92 3.43 40.83 60.40 76.51
0.08 5.91 3.44 41.10 60.61 76.60
0.07 5.93 3.39 41.58 60.80 76.72
0.06 5.79 3.38 41.85 61.01 76.91
0.05 5.77 3.32 42.08 61.07 76.95
0.04 5.79 3.35 42.00 61.13 76.94
0.03 5.73 3.33 42.46 61.50 77.22
0.02 5.81 3.33 42.12 61.01 76.80
0.01 5.86 3.41 41.12 60.56 76.56
0.005 5.96 3.51 40.64 60.12 76.19
0.001 6.20 3.53 39.95 59.08 75.52
Table 5: Evaluation of the effects of the scale of the $\delta$ shift on the St. Peter’s Square scene from Phototourism [20].
Validation of $\delta$ We tested our $\delta$-based solvers for different values of $\delta$ and measured their performance. In general, there is no common value of the $\delta$ shift that leads to the best results on all datasets. This is expected since the precision of the mean-point correspondence depends on many different factors, e.g., the viewing angles of the cameras, the type of the motion, the depth and spatial distributions of the 3D points, etc. We set the value for $\delta$ and the total number of refinement iterations by evaluating their effects on St. Peter’s Square scene from the PhotoTourism dataset [20] which we used for validation and did not include it in other results for PhotoTourism in the paper. Tab. 5 shows how the different settings of the scale of the $\delta$ shift affect the accuracy of the $\delta$-based solvers. Based on these experiments we use $\delta=0.04$ as it provides the best results for 4p3v(M±$\pm\delta$) solver and is also close to the optimal value for its variants. The choice of optimal $\delta$ parameter may be scene-dependent and could potentially be set by using learning-based approaches.
Figure 9: Evaluation of the effects of the number of inner refinement (+R) iterations within 4p3v(M)+R+F solver on St. Peter’s Square scene from PhotoTourism [20]. Shown is the speed-accuracy evaluation with different number of Poselib RANSAC iterations.
Figure 10: Noise experiment showing the pose error measured as $\text{max}\left(0.5(\mathtt{R}_{err}^{12}+\mathtt{R}_{err}^{13}),0.5(\mathbf{t}_{err}^{12}+\mathbf{t}_{err}^{13})\right)$, for the calibrated 4p3v problem (left) and the partial calibrated 4p3vf problem (right) as a function of the noise scale in pixels. Here, $\mathtt{R}_{ij}$ and $\mathtt{t}_{ij}$ are the relative rotation and translation of the $i^{\text{th}}$ and $j^{\text{th}}$ views, respectively.
Inner refinement validation.  We also perform validation of the total number of LM steps in the inner refinement (+R) shown in Fig. 9. We chose the value of 2 for other experiments as it provides the best speed-accuracy trade-off across a range of RANSAC iterations. However, we note that other settings may have very similar performance.
Figure 11: Noise experiment showing $\mathtt{R}_{err}^{12}$ (left) and $\mathbf{t}_{err}^{12}$ (right) as functions of the noise scale in pixels for the calibrated 4p3v problem (top) and the partially calibrated 4p3vf problem (bottom). Here $\mathtt{R}_{12}$ and $\mathtt{t}_{12}$ are the relative rotation and translation between the first two views.
Phototourism [20]
Estimator AVG $(^{\circ})$downarrow MED $(^{\circ})$downarrow AUC@5 $\uparrow$ @10 $\uparrow$ @20 $\uparrow$ Runtime (ms) $\downarrow$
4p3v(HC) [19] 11.41 3.89 37.90 53.76 68.08 176.45
5pt+P3P 19.88 3.41 41.36 57.36 71.13 105.50
4p3v(M) 11.65 4.21 35.38 51.73 66.77 176.77
4p3v(M)+R 10.52 3.47 41.01 56.62 70.25 174.94
4p3v(M)+R+F 10.79 3.58 40.09 55.73 69.56 130.54
4p3v(M±$\pm\delta$) 10.44 3.72 38.82 55.07 69.51 172.06
4p3v(M±$\pm\delta$)+R 19.84 3.27 42.62 58.30 71.74 189.21
4p3v(M±$\pm\delta$)+R+F 10.14 3.40 41.45 57.24 70.92 175.78
4p3v(L) 11.56 4.21 35.32 51.77 66.84 376.31
4p3v(L)+R 10.48 3.47 41.02 56.60 70.25 297.88
4p3v(L±$\pm\delta$init)+R 19.86 3.28 42.60 58.26 71.71 730.90
4p3v(O) 19.64 3.29 42.60 58.54 72.09 158.22
4p3v(O)+R 19.50 3.13 44.01 59.67 72.81 186.30
4p3v(O)+R+F 19.45 3.16 43.82 59.55 72.76 136.36
Cambridge Landmarks [21]
Estimator AVG $(^{\circ})$downarrow MED $(^{\circ})$downarrow AUC@5 $\uparrow$ @10 $\uparrow$ @20 $\uparrow$ Runtime (ms) $\downarrow$
4p3v(HC) [19] 15.12 5.51 24.50 43.42 61.10 164.49
5pt+P3P 13.38 5.17 25.94 45.38 63.16 148.53
4p3v(M) 15.31 5.68 22.93 42.06 60.26 138.33
4p3v(M)+R 14.28 5.29 25.37 44.57 62.33 134.84
4p3v(M)+R+F 14.60 5.39 24.89 43.97 61.71 116.51
4p3v(M±$\pm\delta$) 14.18 5.36 24.72 44.12 62.15 181.21
4p3v(M±$\pm\delta$)+R 13.67 5.11 26.17 45.56 63.33 184.98
4p3v(M±$\pm\delta$)+R+F 13.89 5.23 25.43 44.78 62.70 138.45
4p3v(L) 15.24 5.72 22.98 42.15 60.35 232.13
4p3v(L)+R 14.23 5.27 25.33 44.49 62.30 177.06
4p3v(L±$\pm\delta$init)+R 13.63 5.10 26.11 45.58 63.35 395.43
4p3v(O) 13.85 5.22 25.61 45.19 63.11 130.43
4p3v(O)+R 13.51 5.07 26.60 46.05 63.76 134.68
[…]